Home‎ > ‎Reports‎ > ‎Crew101‎ > ‎

day06


Crew 101 Log Book for March 17 2011

Photos of the day





Commander's Check-In Report

Tom Caillouet Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: EVA 6: 1223 - 1350

Number Of reports: 5 (Commander, Journalist, Engineering, Science, 1 EVA, Chef)

Brief description of Field Mission Results: Successfully gathered samples at one site. Rain forced early end to EVA.

Plans for Tomorrow: Exploration EVA in the morning. Depending on dampness of soil, afternoon EVA will be sample gathering, otherwise another exploration EVA.


Commander's Report

Tom Caillouet Reporting

It has been another successful day on Mars. After a full morning of preparation, Jon and Mitesh took out the ATVs for sample collection. This was also our first chance to do range testing of our radios. While they successfully gathered samples, imaged the test site, and confirmed radio operation, the weather turned against our plans, and it began to rain. This naturally forced us to cut the EVA short, and it also scrubbed the second EVA planned for the day.

With less field activity comes more in hab activity, though. We have spent a large portion of the day preparing a St. Patrick's Day meal of shepard's pie and soda bread and celebrating the holiday with a continuous playlist of Flogging Molly and The Dropkick Murphies. We have also used this extra time to prepare for our NASA INSPIRE presentation later tonight and to prepare for EVAs tomorrow. Hopefully the rain will depart quickly so we can return to full activity soon.


Engineering Report

Cody Hall Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

o V everstart

o V armor plate

o V green


Notes/Comments (include how many generator runs, approximately when,

and any times you needed to turn the generator on/off):

1 generator run


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.75

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.25

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

Honey Oil Quantity (# of Quarts in storage): Leave it to DG

ATV Oil Quantity (# of Quarts in storage): 6 quarts on the

engineering work bench

Car Oil Quantity (# of Quarts in storage): Leave it to DG


Notes/Comments:


Water Status:

Outside Potable Water Tank Level (inches from bottom): 9

Trailer Potable Water Tank Level (inches from bottom): 19

Hab Potable Water Tank Level (inches from bottom): 5

Potable Water Meter Reading: 436128

Notes/Comments :


GreenHab:

Condition of plants in Tank 1: unchanged

Condition of plants in Tank 2: more duckweed

Condition of plants in Tank 3: unchanged

Greenhab notes/ comments:

Internet still bad. Cannot attach pictures.


Observatory:

Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): No

Oil Checked (yes/no): No

Tire Status: Not checked

Odometer at end of day: 123688.4

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: yes

Oil Checked: no

Fuel Consumed: small amount, not replaced

Tire Status: ok

Comments:


ATV 2 (Viking I):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: ok

Comments:


ATV 3 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: ok

Comments


ATV 4 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: small amount, not replaced

Tire Status: ok

Comments:

ATV Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 68

Actual temperature upstairs (Farenheit): 73

Thermostat setting downstairs (Farenheit): 64

Actual temperature downstairs (Farenheit): 63

Notes/Comments:


Computers/Networking Infrastructure: working

Notes/comments on internet/computers:


General Engineering

Notes/Comments:


Science Report

Christina Graves Reporting

On 3-17-11, an ATV EVA (EVA #6) was performed by Mitesh Agrawal and Jonathan Kosh.  The purpose of the EVA was to collect soil samples and geological data from a site on the San Rafael Swell (the site of the MDRS) for analysis.  The San Rafael Swell is characterized by both anticline and sedimentary rocks.  Geological data obtained includes rock morphology, type, and location, as well as Geiger counter readings.

Data obtained from the EVA is found below:

Site 1

Local time: 13:17

GPS Location Data (UTM):  12S 0519191   4250809

Acc (m): 7

Alt (m): 1368

# of rocks larger than 6 in: 9

depth of collection: 0.5 in.

Geiger counter (mR/hr): 0.05-0.1; 0.2 (?)

Crew members Kosh and Agrawal arrived at the site of interest (Site 1) at 13:57. A 6' x 6' area was marked and photographed at 12S 0519191  4250809 UTM (+/- 7m) at an altitude of 1368m (+/- 7m).   A 320.8g soil sample was collected at a depth of 0.5 in. In addition, a Geiger counter reading was taken of the area, with most readings between 0.05 - 0.1 mR/hr. It should be noted that there was one "blip" in the readings which registered at 0.2 mR/hr; however, it is unclear whether this was an accurate reading or a technical artifact.  The area was a water run-off area; chosen because of the probability of bacterial variety at the site.  The site also serves as a potential model of a "primordial pool," thought to be necessary for life in prebiotic situations.

The sample collected is a uniform sample of fine sand. There were several (9) larger rocks in the area, although none were collected for purposes of this experiment.  Larger rocks were classified as those greater than 6 in. in diameter.

Sample collection was limited on this EVA due to inclement weather (rain); therefore, samples were only obtained from 1 site.

Biological soil analysis (EZ N.A. Soil Analysis from Omega and PCR) will be performed at a later time to test for the presence of bacteria.



EVA 6 Report

Jon Kosh Reporting

EVA Time: 1223 - 1350

EVA Crewmembers:

Jon (KF7GLD)

Mitesh (KK4AQI)

Base Communications: Tommy (KJ4KAC)

EVA Procedure: Sample soil at sites where water flow characteristics are evident. General site locations north, northeast of the Hab were selected beforehand based on satellite images. EVA team not to deviate from established paths while on ATVs.  Before exploring on foot, ATV helmets are to be exchanged with bubble helmets.  Once the sample site is reached, stakes are to mark out a 6' x 6' area. Pictures of the test site will be taken from each stake, then a sample will be gathered from the area. Radiation readings and a count of large (>6" characteristic length) rocks will also be taken. This will be performed at up to five sites.

EVA Results: Soil sample returned from a promising site located at 12N 0519191, 4251413 (UTM). After evaluating weather conditions, the team returned to Hab, safely inside from the Red Planet rain.

Lessons Learned: Start EVAs earlier in the day, stay mindful of weather conditions.


Chef's Report

Christina Graves Reporting

1a) Today was a cooking day.

1b) Was today a special day celebrated at the main meal with special

food or activity?

Yes - St. Patrick's Day!

2) List the foods served at today's main meal (usually the evening

meal), giving full names of each.

Shepherd's Pie

3) List any main meal foods not finished by the crew (leftovers stored

for later use or discarded)

none

4) Recipes for anything more complicated than following package

directions exactly, or rehydrating.

Potatoes

1 1/2 c. diced potatoes

1/2 c. milk

1tbl. butter

salt and pepper to taste

2 heaping tbl. mayonnaise

Beef and Vegetables

1 c. dehydrated beef (you can use a different type of meat, if you prefer)

1 cube beef or vegetable bouillon

1/4 c. celery

1/4 c. carrots

1/4 c. onion

1 c. peas

1 tsp. pepper

1 tsp. oil or crisco

1/4 tsp. salt (or to taste)

1/4 tsp. dill

Rehydration - do this 2-3 hours prior to cooking

In separate containers, rehydrate:

1. beef with bouillon

2. peas, onions, carrots, celery

3. diced potatoes

Cooking beef and vegetables

1.  heat oil or crisco in pan

2. add beef with bouillon; heat until thoroughly cooked

3. Season with pepper, dill, and salt to taste

4. add vegetables and cook again until thoroughly heated

Potatoes

In a separate pot:

1. add water to rehydrated potatoes, and hard boil until very soft

2. drain

3. add mayonnaise and milk, mashing with a fork (you can also use a

hand mixer if available)

4. add salt and pepper to taste

In a baking dish, layer bottom of dish with meat and vegetables. Cover the meat and vegetable layer with potatoes. Bake at 350F (in oven) until thoroughly cooked. Top mashed potatoes with melted butter to brown.

5) Lessons learned (if any)

ketchup would be a good condiment with this dish.

6) Comments/questions for Kim and Jean, the food study investigators

None


Journalist Report

Christine Redmond Reporting

Happy Saint Patty’s day earthlings! It has been a great day here on Mars! This morning we mapped out our 6th EVA on an aerial map of the site and Jon and Mitesh headed out to execute this EVA shortly afterward. The ATVs are back up and running so they were able to gather samples for our geology and biology experiments from a site further from home base. Tommy stayed back and manned the radio station for this EVA.  The EVA team had to cut their EVA slightly short due to weather. I was able to collect some sufficient dirt and dust for our contamination study upon their return to base.

This morning I made soda bread for our St. Patty’s day dinner tonight and Christina has been busy making a Sheppard’s pie this afternoon. We are all ready to celebrate St. Patty’s day from Mars! But before the celebration begins we have one more official task for the day, our NASA INSPIRE outreach chat! We have all been looking forward to sharing our experience with the INSPIRE students and answering the great questions they are sure to have for us.